A Parent Company owns 100% of its Subsidiary. During 2013, the Parent company reports net income of $1,000,000 and the subsidiary reports net income of $400,000. The Parent had a bond payable outstanding on July 1, 2012, with a carry value equal to $840,000. The Subsidiary acquired the bond on July 1, 2012 for $790,000. During 2013, the Parent reported interest expense (related to the bond) of $70,000 while the Subsidiary reported interest income (related to the bond) of $64,000. What is consolidated net income for the year ended December 31, 2013?

a.   $1,400,000

b.   $1,406,000

c.   $1,450,000

d.   $1,456,000


Please provide step by step explanation for the answer.
Answer #1

Parent company Income = 1,000,000

Subsidairy Income = 400,000

Consolidated Income = 1,000,000 + 400,000 + 6000 = 1,406,000 ................ Option - B

cancellation of excess of Interest expenses over Income = 70000 - 64000 = 6,000